Four times the sum of a number and 5 equals 64.
Answer by sdmmadam@yahoo.com(530)

Four times the sum of a number and 5 equals 64.
To find the number.
Let the number be n
Four times the sum of n and 5 equals 64.
That is 4X(n+5) = 64
4n +20 = 64
4n = 64-20
4n = 44
Dividing by 4
n = 44/4 = 11
Answer: The required number is 11
Verification: Four times the sum of n and 5 should be equal to 64.
4X(11+5) = 4X16 = 64 which is right